QUESTION 1
(12 marks; 22 minutes)


Poullus Limited (PL). a high-tech electronics manufacturing company, is currently operating at 70% of its maximum capacity A new customer in lesotho has asked PL to provide a special order for 200 virtual reality headsets at R2.500 per unit. The potential customer is not a current customer of PL, but the directors of PL are keen to try and win the contract as they believe that this may lead to more contracts in the future from Lesotho. As a result, they intend pricing the contract using relevant costs.

The following information has been obtained from a four-hour meeting that the Production Manager of PL had with the potential customer:
1) .400 kilograms of material A will be required. This is a material that is regularly used by PL and there are 700 kilograms currently in inventory. These were bought at a cost of R250 per kilogram. They have a resale value of R220 per kilogram and their current replacement cost is R260 per kilogram.

2). 800 kilograms of material B will be required. This material will have to be purchased for the contract because it is not otherwise used by PL. The minimum order quantity
from the supplier is 1 000 kilograms at a cost of R180 per kilogram. PL does not expect to have any use for any of this material that remains after this contract is completed.

3) 200 components will be required. These will be purchased from Xboks Limited. The purchase price is R1 000 per component.

4) A total of 260 direct labour hours will be required. The current wage rate for the appropriate grade of direct labour is R280 per hour. Currently PL has 80 direct labour hours of spare capacity at this grade that is being paid under a guaranteed wage agreement. The additional hours would need to be obtained by either (1) overtime at a total cost of R340 per hour; or (ii) recruiting temporary staff at a cost of R300 per hour. However, if temporary staff are used they will not be as experienced as PL's existing workers and will require 10 hours supervision by an existing supervisor who would be paid overtime at a cost of R420 per hour for this work.

5) 20 machine hours will be required. The machine to be used is already leased for a weekly leasing cost of R 12,000. It has a capacity of 40 hours per week. The machine has sufficient available capacity for the contract to be completed. The variable running cost of the machine is R140 per hour.

The company absorbs its fixed overhead costs using an absorption rate of R400 per direct labour hour.

The Production Manager is paid an annual salary equivalent to R15,800 per 8-hour day. Pl's practice is to establish selling prices by applying a mark-up on cost of 45%.

Required:

Calculate the relevant cost of building 200 virtual headsets which may help with the pricing decision.

NOTE: Present your answer in a schedule that clearly shows the relevant cost value for each of the items identified above and explain each relevant cost value you have included in your schedule and why the values you have excluded are not relevant
Answer #1

Poullus Limited

Statement Showing Relevant cost of manufacturing 200 Virtual headsets:

Particulars Amount Amount (R) per unit
Material Cost (Note-1) 484000 2420
Labour (Note-2) 58200 291
Variable Machine running cost (Note-3) 2800 14
Total relevant cost for 200 virtual headsets

545000

2725

Statement showing Total cost for the purpose of Pricing decision:

Particulars Working Amount (R)
Total Relevant cost 545000
Labour cost for spare capacity 80*280 22400
Cost of Leasing Machine (Note 4) 12000*20/40 6000
Fixed Overhead Cost R400*260 hrs 104000
Production Managers Salary 15800*260/8 513500
Total cost for 200 virtual headsets 11,90,900

The relevant cost of producing virtual headsets i.e. R 2725 per unit exceeds the quoted price of R 2500 per unit. Hence, it is not viable for the company to accept the offer.

Note-1: Material Cost:

a) Material A: Requirement= 400 kilograms

Since, Material A is regularly used, any requirement for the special order shall have to be replaced. Hence, cost of purchase of 700 kgs of Material A which is already lying in the stock is not relevant for the pricing decision. Also, the Resale value of R 220 per kg is of no use in the pricing decision.

Hence, relevant cost of Material A is the Current purchase price of the material i.e. 400*260 = R 104000

b) Material B: Requirement= 800 kgs

Material B is not regularly used by PL and there is no stock of B lying with PL. Further, the material can only be purchased with a minimum order quantity (MOQ) of 1000kgs. So, we shall purchase the entire MOQ of 1000kgs and charge the full purchase price to the order, as there will be no alternatie use of this material in the future.

Hence, relevant cost Material B= 1000*180= R180000.

c) Component: Requirement= 200 units

Relevant cost= 200*1000 = R 200000

Hence, Total Material cost= a+b+c = R(104000+180000+200000) = R 484000

Note 2: Labour Cost-

Total Direct Labour hours required= 260 hrs.

The company currently has a spare capacity of 80 hrs and the labour is being paid under a guaranteed wage agreement. Hence, any payment made upto 80 hrs is not relevant for decision making as the cost will be incurred irrespective of whether or not the order is accepted.

For the remaining 180 hrs we have two options. We shall select the option which has lower total direct labour cost. Let us now evaluate the two options-

Option 1- Overtime @340 per hr-

Total Direct Labour Cost- 180*340= R 61200

Option 2: Recruit Temporary Staff

Wages payable to temporary staff 180*300 R 54000
Supervisors Overtime Charge 10*420 R 4200
Total R 58200

Hence, we select option 2 and total cost of direct labour= R 58200.

Note 3: Variable Machine running cost-

Total machine hrs required= 20 hrs

Variable Machine running cost= 20*140 = R 2800.

Note 4: Cost of leasing-

The machine has already been taken on lease and so the Cost of leasing R 12000/- is not relevant if a decision is to be made regarding acceptance of order. However, it is relevant for pricing decision. Also, the machine is already leased and has a total capacity of 40 hrs. We have assumed that the machine is used for other purposes as well. Hence, we have considered only half cost of leasing in total cost.